# Cross-Service Threat Intelligence in LLM Services using Privacy-Preserving Fingerprints

### v1

@@ 2.1.3 BinaryShield’s Randomized-Response Layer
Randomized response inevitably perturbs the underlying signal. The expected self-distance introduced by the mechanism is
@@
$\mathrm{H}(\mathbf{b},\mathbf{\tilde{b}})=\sum_{i=1}^{d}\mathbb{I}[b_{i}\neq\tilde{b_{i}}$]
@@
where $\mathbb{I}[\cdot]$ is the indicator function, which equals 1 if its argument is true and 0 otherwise. Because $p$ remains close to one even for moderate $\alpha$, the added noise is sparse and, as we show experimentally in Section 3, preserves enough structure for accurate approximate matching while providing the formal privacy guarantees demanded by cross-service threat intelligence.
@@
The randomization budget $\alpha$ serves as a control parameter governing the privacy-utility tradeoff in BinaryShield. When $\alpha$ approaches zero, the mechanism maximizes privacy protection (i.e., infinite privacy) but renders fingerprints unusable for correlation. As $\alpha$ increases, utility improves while maintaining privacy guarantees in a controlled manner. Figure 4 illustrates this relationship through PCA visualization of binary vectors before and after applying the Randomized Response mechanism (Warner, 1965). At $\alpha=0.25$, privatized vectors (red triangles) appear completely random. As $\alpha$ increases (reducing noise), these vectors progressively align closer to their original non-private counterparts (blue circles). Notably, even at $\alpha=3$, a non-zero average displacement between original and privatized vectors persists, making reconstruction of the original prompt challenging. This visualization confirms that larger $\alpha$ values preserve more of the vector’s structure while maintaining formal differential privacy guarantees through controlled bit-flipping. In practice, higher privacy budgets may be appropriate for intra-service sharing where additional security measures already exist within organizational boundaries.
2.1.4 BinaryShield’s Cross-Service Threat Correlation
@@
When a service detects a potential prompt injection, it generates a fingerprint using the BinaryShield pipeline. For clarity, we denote the locally randomized-response fingerprint $\tilde{\mathbf{b}}\in\{0,1\}^{d}$ as $f$ when broadcasting and correlating across services (i.e., $f\approx\tilde{\mathbf{b}}$). This fingerprint is broadcast asynchronously to participating services through secure channels. Upon receiving a fingerprint, services scan their historical logs for similar instances. If matches are found, alerts are triggered and defenses can be updated proactively. To maintain privacy and compliance, only aggregate match statistics are shared, and no specific prompt content is revealed. This workflow enables rapid propagation of threat intelligence while respecting organizational boundaries. Formally, let $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{N}\}$ denote the set of $N$ services, each with its own compliance boundary. Each service $S_{i}$ maintains a log of previously observed fingerprints $\mathcal{F}_{i}=\{f_{i}^{(1)},f_{i}^{(2)},\ldots,f_{i}^{(M_{i})}\}$, where $f_{i}^{(m)}\in\{0,1\}^{d}$ is a $d$-dimensional binary fingerprint. Suppose service $S_{q}$ detects a suspicious prompt and generates a fingerprint $f_{q}\in\{0,1\}^{d}$ using the BinaryShield pipeline. This fingerprint is securely broadcast to all other services $S_{j}$ ($j\neq q$). Each recipient service $S_{j}$ performs the following search:
$\mathcal{M}_{j}(f_{q})=\left\{f_{j}^{(m)}\in\mathcal{F}_{j}\;\middle|\;\mathrm{H}(f_{q},f_{j}^{(m)})\leq\tau\right\}$
@@
where $\mathcal{M}_{j}(f_{q})$ is the set of fingerprints belonging to service $S_{j}$ (i.e., fingerprints in $\mathcal{F}_{j}$) that match $f_{q}$ within a Hamming distance threshold $\tau$, and $\mathrm{H}(\cdot,\cdot)$ is the Hamming distance as previously defined. This protocol enables rapid, privacy-preserving propagation of threat intelligence across all $N$ services. This protocol generalizes to any number of services. For example, if Service 2 ($S_{2}$) generates a fingerprint $f_{2}$, it broadcasts $f_{2}$ to Services 1 and 3 ($S_{1}$, $S_{3}$), which independently search their logs for similar fingerprints. If $S_{1}$ finds two matches and $S_{3}$ finds one, these services take action based on their policy and only the counts are reported back, enabling rapid, privacy-preserving propagation of threat intelligence across organizational boundaries.
Overall, when a match is found below a calibrated threshold, the service can take protective actions such as updating classifier parameters, flagging suspicious activity, or blocking potential attacks. Services can subscribe to threat intelligence feeds from peer services, automatically incorporating new attack patterns into their defense mechanisms.
Summary.
@@
BinaryShield offers a scalable and privacy-preserving solution for sharing prompt injection threat intelligence across organizational and compliance boundaries. By combining PII redaction, semantic embedding, binary quantization, and differential privacy, the system achieves strong privacy protections without sacrificing detection performance or scalability. The modular architecture allows for privacy parameters to be tuned to specific requirements, while efficient processing supports production-scale deployments. As large language model adoption expands, BinaryShield establishes a foundation for collaborative and adaptive defenses against evolving prompt injection attacks.
3 Evaluation
We conduct a comprehensive evaluation of BinaryShield to assess its effectiveness in privacy-preserving threat intelligence sharing for prompt injection attacks. Our evaluation addresses key research questions across diverse, systematically generated attack variants. BinaryShield’s evaluation is structured around the following research questions:
@@ 3 Evaluation
  • RQ1 (Detection Effectiveness). How accurately does BinaryShield detect prompt injection attacks across systematically generated adversarial variants, from simple word substitutions to comprehensive semantic paraphrasing, compared to the SimHash baseline?
@@
  • RQ5 (Computational Efficiency). What are the computational and storage efficiency gains of BinaryShield’s binary fingerprints compared to dense embeddings, and do these improvements enable practical real-time threat correlation at enterprise scale?
3.1 Evaluation Settings
3.1.1 Datasets
@@
We evaluate BinaryShield using the recent prompt injection dataset from (Shen et al., 2024), which contains real-world attack and benign prompts. From this, we construct a comprehensive evaluation set simulating adversarial scenarios, enabling controlled and realistic assessment of threat intelligence in LLM services.
@@
Given the inherent challenge of obtaining large-scale labeled datasets of semantically similar prompt injection variants, we develop a systematic methodology for generating controlled attack modifications to validate threat intelligence. Our synthetic data generation process creates two primary types of variants: Word-Flipping Variants: For each adversarial prompt, we randomly select $x$ words (where $x\in\{1,3,5,10,20\}$) longer than four characters and replace them with semantically equivalent synonyms using GPT-4o. This simulates subtle adversarial modifications that maintain attack intent while altering surface-level tokens. The language model is guided using the structured prompt: "Understand this adversarial prompt: <original prompt>. Now generate synonyms for the following words: <words to change>. Return the JSON with the field ’words_changes_to’ containing the synonyms in order." Paraphrase Variants: We generate comprehensive paraphrased versions of prompts using GPT-4o while preserving the underlying attack intent and malicious purpose. The generation prompt instructs the model: "Paraphrase the following adversarial prompt while maintaining the attack intent and purpose. Original prompt: <original prompt>. Return the JSON with the field ’generated_response’." Benign Pairs: Benign prompts are systematically paired such that each pair consists of two semantically unrelated prompts, ensuring they serve as appropriate negative examples for similarity detection algorithms (e.g., SimHash and BinaryShield). The resulting dataset undergoes rigorous filtering to ensure quality and consistency (e.g., remove duplicates, empty prompts). The final dataset is balanced with equal numbers of attack and benign pairs, with attack pairs labeled as 1 and benign pairs labeled as 0. This balanced approach ensures unbiased evaluation across both positive and negative cases. Our synthetic generation methodology enables systematic evaluation across varying levels of attack sophistication, from minimal single-word modifications to comprehensive paraphrasing that significantly alters surface structure while preserving malicious intent. This controlled approach allows us to assess algorithm robustness across the spectrum of adversarial modifications likely encountered in real-world deployment scenarios.
3.1.2 Baseline
Our evaluation compares two primary approaches for prompt similarity detection BinaryShield and SimHash. BinaryShield utilizes ModernBert  for semantic feature extraction, followed by sign-based binary quantization and randomized response for local differential privacy. We also evaluate it with OpenAI embedding model to demonstrate the flexibility of BinaryShield in supporting both open source and closed source models. SimHash is a widely-adopted locality-sensitive hashing algorithm that serves as our primary baseline. SimHash generates compact binary fingerprints and has been extensively used for near-duplicate detection in large-scale text corpora (Manku et al., 2007).
3.1.3 Evaluation Metrics
@@
We employ a comprehensive set of metrics to assess both detection performance and privacy-utility tradeoffs. For detection performance, we use precision, recall, accuracy, and F1-score across varying similarity thresholds to evaluate attack detection accuracy. Threshold analysis is performed using precision-recall curves to identify the optimal threshold for Hamming distance and to understand how performance is affected by threshold selection, as discussed in (Manku et al., 2007). We also include confusion matrices for detailed analysis of true positives (TP), false positives (FP), true negatives (TN), and false negatives (FN) at optimal thresholds.
@@
Figure 6: SimHash baseline performance analysis across attack variants. Confusion matrices reveal growing FP and FN, particularly for paraphrasing scenarios where token-based similarity detection struggles.
3.2 BinaryShield Comparison with Baseline
To provide a rigorous performance analysis, we systematically compare BinaryShield against SimHash, a widely used locality-sensitive semantic hashing method (Manku et al., 2007). Both algorithms generate binary fingerprints and use Hamming distance for similarity measurement, making them directly comparable. This baseline is chosen to evaluate not only detection performance but also the privacy-preserving capabilities essential for cross-boundary threat intelligence sharing as SimHash also generates non-reversible binary fingerprints.
For this evaluation, we configure BinaryShield with a differential privacy parameter of $\alpha=2.0$, which provides meaningful privacy protection while maintaining strong detection performance. We note that BinaryShield with ModernBert generates 768-dimensional binary vector while SimHash produces 64-dimensional vector.
Our comparative analysis encompasses six distinct attack variants: minimal single-word modifications (V-1), progressive word changes (V-5, V-20) and extensive modifications in paraphrasing scenarios (V-Paraphrase) that fundamentally restructure prompts while preserving malicious intent. This progression allows us to assess how each method’s performance degrades as attack sophistication increases during threat correlation. Figure 5 and Figure 6 present comprehensive performance analysis through precision-recall curves (subfigures -a) and confusion matrices at optimal operating points (subfigures -b).
Figure 7: F1-score comparison between BinaryShield and SimHash across attack variants. BinaryShield consistently outperforms SimHash, particularly in complex paraphrasing scenarios, demonstrating superior robustness to semantic modifications.
@@ 3.2 BinaryShield Comparison with Baseline
As attack complexity increases to extensive word modifications (V-20), performance differences begin to emerge. BinaryShield maintains robust performance with an F1-score of 0.99 (precision: 0.99, recall: 0.98), while SimHash experiences measurable degradation to an F1-score of 0.96 (precision: 0.98, recall: 0.93). The 3.3 percentage point gap starts to illustrate BinaryShield’s superior stability as lexical modifications increase.
@@
The most challenging scenario involves comprehensive paraphrasing attacks (V-Paraphrase), where prompts undergo fundamental restructuring while preserving malicious intent. Here, the performance differential becomes most pronounced: BinaryShield achieves an F1-score of 0.94 (precision: 0.92, recall: 0.95), while SimHash degrades significantly to an F1-score of 0.77 (precision: 0.75, recall: 0.80). This 17 percentage point gap in F1 score demonstrates BinaryShield’s fundamental advantage in threat intelligence that transcends surface-level linguistic variations.
@@
Additionally, examining confusion matrices at optimal thresholds, provides additional insights into error patterns in Figures  5(b) and  6(b). BinaryShield demonstrates consistent performance as compared to SimHash. SimHash shows increasing error variability as attack complexity increases, struggling particularly to distinguish between semantically similar malicious content and unrelated benign prompts in paraphrasing scenarios.
Our empirical results demonstrate that BinaryShield provides superior robustness for prompt injection threat correlation compared to SimHash, with advantages becoming most pronounced for sophisticated attacks that represent complicated adversarial scenarios. The 17 percentage point F1-score gap in paraphrasing attacks, the most challenging and operationally relevant scenario, validates the effectiveness of BinaryShield over semantic hashing technique like SimHash. These performance improvements are achieved while simultaneously providing privacy guarantees essential for cross-boundary threat intelligence sharing. The superior performance on complex attacks is particularly significant, as sophisticated paraphrasing and extensive modification represent the most realistic attack patterns likely to be encountered in production environments where adversaries actively attempt to evade detection systems.
@@
Summary. BinaryShield consistently outperforms SimHash in prompt injection threat correlation, especially under complex paraphrasing attacks, achieving an F1-score of 0.94 compared to SimHash’s 0.77. These results highlight BinaryShield’s superior robustness with enabling privacy, making it highly effective for cross-boundary threat intelligence sharing.
@@
Figure 8: Privacy-utility trade-off for paraphrased prompt injection detection. (Left) BinaryShield F1 rises smoothly with privacy budget $\alpha$, dominating SimHash (privacy-preserving baseline) across most of the spectrum and approaching non-private dense embeddings. (Center) Precision lags recall at high noise, then rapidly converges after $\alpha\!\approx\!1$ as false positives collapse. Horizontal dashed lines: SimHash precision/recall. (Right) Confusion matrices at extreme settings: severe noise ($\alpha=0.25$) induces many false positives; higher $\alpha$ (33) sharply improves precision.
3.2.1 Empirical Analysis of Randomized Response Impact

### v2

@@ II-A3 BinaryShield’s Randomized-Response Layer
This DP mechanism inevitably perturbs the fingerprint. The expected self-distance introduced by it is
@@
$\mathrm{H}(\mathbf{b},\mathbf{\tilde{b}})=\sum_{i=1}^{d}\mathbb{I}[b_{i}\neq\tilde{b_{i}}$],
@@
where $\mathbb{I}[\cdot]$ is the indicator function, which equals 1 if its argument is true and 0 otherwise. Because $p$ remains close to one even for moderate $\alpha$, the added noise is sparse and, as we show experimentally in Section III, preserves enough structure for accurate approximate matching while providing the privacy guarantees demanded by cross-service threat intelligence.
@@
The randomization budget $\alpha$ serves as a control parameter governing the privacy-utility tradeoff in BinaryShield. When $\alpha$ approaches zero, the mechanism maximizes privacy protection (i.e., infinite privacy) but renders fingerprints unusable for correlation. As $\alpha$ increases, utility improves while maintaining privacy guarantees in a controlled manner. Figure 4 illustrates this relationship through PCA visualization of binary vectors before and after applying the Randomized Response mechanism [51]. At $\alpha=0.25$, privatized vectors (red triangles) appear completely random. As $\alpha$ increases (reducing noise), these vectors progressively align closer to their original non-private counterparts (blue circles). Notably, even at $\alpha=3$, a non-zero average displacement between original and privatized vectors persists, making reconstruction of the original prompt challenging. This visualization confirms that larger $\alpha$ values preserve more of the vector’s structure while maintaining formal differential privacy guarantees through controlled bit-flipping. In practice, higher privacy budgets may be appropriate for intra-service sharing where additional security measures already exist within organizational boundaries.
II-A4 BinaryShield’s Cross-Service Threat Correlation
@@
When a service detects a potential prompt injection, it generates a fingerprint using the BinaryShield’s pipeline. For clarity, we denote the locally randomized-response fingerprint $\tilde{\mathbf{b}}\in\{0,1\}^{d}$ as $f$ when broadcasting and correlating across services (i.e., $f\approx\tilde{\mathbf{b}}$). This fingerprint is broadcast asynchronously to participating services. Upon receiving a fingerprint, services scan their historical logs for similar instances. If matches are found, alerts are triggered and defenses can be updated proactively. To maintain privacy and compliance, only aggregate match statistics are shared, and no specific prompt content is revealed. This workflow enables rapid propagation of threat intelligence while respecting organizational boundaries. Formally, let $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{N}\}$ denote the set of $N$ services, each with its own compliance boundary. Each service $S_{i}$ maintains a log of previously observed fingerprints $\mathcal{F}_{i}=\{f_{i}^{(1)},f_{i}^{(2)},\ldots,f_{i}^{(M_{i})}\}$, where $f_{i}^{(m)}\in\{0,1\}^{d}$ is a $d$-dimensional binary fingerprint. Suppose service $S_{q}$ detects a suspicious prompt and generates a fingerprint $f_{q}\in\{0,1\}^{d}$ using the BinaryShield pipeline. This fingerprint is securely broadcast to all other services $S_{j}$ ($j\neq q$). Each recipient service $S_{j}$ performs the following search:
$\mathcal{M}_{j}(f_{q})=\left\{f_{j}^{(m)}\in\mathcal{F}_{j}\;\middle|\;\mathrm{H}(f_{q},f_{j}^{(m)})\leq\tau\right\},$
@@
where $\mathcal{M}_{j}(f_{q})$ is the set of fingerprints belonging to service $S_{j}$ (i.e., fingerprints in $\mathcal{F}_{j}$) that match $f_{q}$ within a Hamming distance threshold $\tau$, and $\mathrm{H}(\cdot,\cdot)$ is the Hamming distance as previously defined. This enables rapid, privacy-preserving propagation of threat intelligence across all $N$ services. This protocol generalizes to any number of services. For example, if Service 2 ($S_{2}$) generates a fingerprint $f_{2}$, it broadcasts $f_{2}$ to Services 1 and 3 ($S_{1}$, $S_{3}$), which independently search their logs for similar fingerprints. If $S_{1}$ finds two matches and $S_{3}$ finds one, these services take action based on their policy and only the counts are reported back, enabling rapid, privacy-preserving propagation of threat intelligence across organizational boundaries.
@@
Summary. BinaryShield offers a scalable and privacy-preserving solution for sharing prompt injection threat intelligence across organizational compliance boundaries. By combining PII redaction, semantic embedding, binary quantization, and differential privacy, the system achieves strong privacy protections without sacrificing detection performance or scalability. The modular architecture allows for privacy parameters to be tuned to specific requirements, while efficient processing supports production-scale deployments. As LLM adoption expands, BinaryShield establishes a foundation for collaborative defenses against evolving attacks.
III Evaluation
We present a comprehensive evaluation of BinaryShield, structured around the following research questions:
@@ III Evaluation
  • RQ1 (Detection Effectiveness). How accurately does BinaryShield match prompt injection attacks across systematically generated adversarial variants, from simple word substitutions to comprehensive semantic paraphrasing, compared to the SimHash [37] baseline?
@@
  • RQ5 (Computational Efficiency). What are the computational gains of BinaryShield’s fingerprints compared to dense embeddings, and do these improvements enable practical real-time threat correlation at enterprise scale?
III-A Evaluation Settings
III-A1 Datasets
@@
We evaluate BinaryShield using the recent prompt injection dataset from [45], which contains real-world attack and benign prompts. From this, we construct a comprehensive evaluation set simulating adversarial scenarios, enabling controlled and realistic assessment of threat intelligence.
@@
Given the inherent challenge of obtaining large-scale labeled datasets of semantically similar prompt injection variants, we develop a systematic methodology for generating controlled attack modifications to validate threat intelligence. Our synthetic data generation process creates two primary types of variants: Word-Flipping Variants: For each adversarial prompt, we randomly select $x$ words (where $x\in\{1,3,5,10,20\}$) longer than four characters and replace them with semantically equivalent synonyms using GPT-4o. This simulates subtle adversarial modifications that maintain attack intent while altering surface-level tokens. The language model is guided using the structured prompt listed in Appendix LABEL:prompt:words-flipped. Paraphrase Variants: We generate comprehensive paraphrased versions of prompts using GPT-4o while preserving the underlying attack intent. Appendix LABEL:prompt:paraphrase contains the prompt. Benign Pairs: Benign prompts are systematically paired such that each pair consists of two semantically unrelated prompts, ensuring they serve as appropriate negative examples for similarity detection algorithms (e.g., SimHash and BinaryShield). The resulting dataset undergoes rigorous filtering to ensure quality and consistency (e.g., remove duplicates, empty prompts). The final dataset is balanced with equal numbers of attack and benign pairs, with attack pairs labeled as 1 and benign pairs labeled as 0. This balanced approach ensures unbiased evaluation across both positive and negative cases. Our synthetic generation methodology enables systematic evaluation across varying levels of attack sophistication, from minimal single-word modifications to comprehensive paraphrasing.
III-A2 Baseline
Our evaluation compares BinaryShield against SimHash [37], a widely-adopted locality-sensitive hashing baseline. BinaryShield operates by first extracting semantic features using an embedding model. To demonstrate its flexibility with both open-source and proprietary options, we evaluate BinaryShield using ModernBert [23] and an OpenAI model.
III-A3 Evaluation Metrics
@@
We employ a comprehensive set of metrics to assess both detection performance and privacy-utility tradeoffs. For detection performance, we use precision, recall, accuracy, and F1-score across varying similarity thresholds to evaluate attack detection accuracy. Threshold analysis is performed using precision-recall curves to identify the optimal threshold for Hamming distance and to understand how performance is affected by threshold selection, as discussed in [37]. We also include confusion matrices for detailed analysis of true positives (TP), false positives (FP), true negatives (TN), and false negatives (FN) at optimal thresholds.
@@
Figure 6: SimHash baseline performance analysis across attack variants. Confusion matrices reveal growing FP and FN, particularly for paraphrasing scenarios.
III-B BinaryShield Comparison with Baseline
To provide a rigorous performance analysis, we systematically compare BinaryShield against SimHash, a widely used locality-sensitive semantic hashing method [37]. Both algorithms generate binary fingerprints and use Hamming distance for similarity measurement, making them directly comparable. This baseline is chosen to evaluate not only detection performance but also the privacy-preserving capabilities essential for cross-boundary threat intelligence sharing as SimHash also generates privacy-preserving binary fingerprints.
For this evaluation, we configure BinaryShield with a differential privacy parameter of $\alpha=2.0$, which provides meaningful privacy protection while maintaining strong detection performance. Our comparative analysis encompasses six distinct attack variants: minimal single-word modifications (V-1), progressive word changes (V-5, V-20) and extensive modifications in paraphrasing scenarios (V-Paraphrase) that fundamentally restructure prompts while preserving malicious intent. This progression allows us to assess how each method’s performance degrades as attack sophistication increases during threat correlation. Figure 5 and Figure 6 present comprehensive performance analysis through precision-recall curves and confusion matrices at optimal operating points.
Figure 7: F1-score comparison between BinaryShield and SimHash across attack variants. BinaryShield outperforms SimHash, particularly in complex paraphrasing scenarios, demonstrating superior robustness to semantic modifications.
@@ III-B BinaryShield Comparison with Baseline
As attack complexity increases to extensive word modifications (V-20), performance differences begin to emerge. BinaryShield maintains robust performance with an F1-score of 0.99 (precision: 0.99, recall: 0.98), while SimHash experiences measurable degradation to an F1-score of 0.96 (precision: 0.98, recall: 0.93). The 3.3% gap starts to illustrate BinaryShield’s superior stability as lexical modifications increase.
@@
The most challenging scenario involves comprehensive paraphrasing attacks (V-Paraphrase), where prompts undergo fundamental restructuring while preserving malicious intent. Here, the performance differential becomes most pronounced: BinaryShield achieves an F1-score of 0.94 (precision: 0.92, recall: 0.95), while SimHash degrades significantly to an F1-score of 0.77 (precision: 0.75, recall: 0.80). This 17% gap in F1 score demonstrates BinaryShield’s fundamental advantage in threat intelligence that transcends surface-level linguistic variations.
@@
Additionally, examining confusion matrices at optimal thresholds, provides additional insights into error patterns in Figures  5(b) and  6(b). BinaryShield demonstrates consistent performance as compared to SimHash. SimHash shows increasing error variability as attack complexity increases, struggling particularly to distinguish between semantically similar malicious content and unrelated benign prompts. The 17 percentage point F1-score gap in paraphrasing attacks, the most challenging and operationally relevant scenario, validates the effectiveness of BinaryShield. These performance improvements are achieved while simultaneously providing privacy guarantees essential for cross-service threat intelligence.
@@
Summary. BinaryShield consistently outperforms SimHash in threat correlation, especially under complex paraphrasing attacks, achieving an F1-score of 0.94 compared to SimHash’s 0.77. These results highlight BinaryShield’s superior robustness with enabling privacy, making it highly effective for cross-service threat intelligence.
@@
Figure 8: Privacy-utility trade-off for paraphrased prompt injection detection. (Left) BinaryShield F1 rises smoothly with privacy budget $\alpha$, dominating SimHash (privacy-preserving baseline) across most of the spectrum and approaching non-private dense embeddings. (Center) Precision lags recall at high noise, then rapidly converges after $\alpha\!\approx\!1$ as FP collapse. Horizontal dashed lines: SimHash precision/recall. (Right) Confusion matrices at extreme settings: severe noise ($\alpha=0.25$) induces many FP; higher $\alpha$ (33) sharply improves precision.
III-B1 Empirical Analysis of Randomized Response Impact